Time Value of Money

 Let us take APR = 6 %. Now calculate the EAR for continuous compounding. 
Answer #1

Annual Percentage Rate, APR = 6% = 0.06

Formula: Effective Annual Rate, EAR = eAPR - 1

∴ EAR = e0.06 - 1

=> EAR = 1.0618 - 1 = 0.0618

Therefore, EAR = 6.18%

answered by: Zahidul Hossain